We read your photo as:

A dartboard has regions A, B, C, and D with the following probabilities of a dart landing in each: P(A) = 0.12, P(B) = 0.18, P(C) = 0.25, P(D) = 0.45. What is the probability that the dart will not la

Mathematics

ScanSolve AI Answer

Step-by-step solution

2 steps

Here are the solutions to the questions.

38. Probability

  1. Identify the given probabilities.

    The probabilities of a dart landing in regions A, B, C, and D are:
    P(A)=0.12P(A) = 0.12
    P(B)=0.18P(B) = 0.18
    P(C)=0.25P(C) = 0.25
    P(D)=0.45P(D) = 0.45

  2. Calculate the probability that the dart will not land in region D.

    The probability of an event not happening is 1 minus the probability of the event happening.
    So, P(not D)=1−P(D)P(\text{not D}) = 1 - P(D).
    Alternatively, $P(\text{

Was this helpful?

Still stuck on this one?

Ask a follow-up, and the answer starts from this question.

Got a different question?Ask your own question
Handwritten step-by-step solution preview

Handwritten Step-by-Step Solution

Get this as a handwritten image on WhatsApp — like a tutor wrote it for you.

Try on WhatsApp